OpenStudy (anonymous):

find w(4) for the function w(x)=3x+7

OpenStudy (anonymous):

X=12

OpenStudy (anonymous):

\[w(x)=3x+7\] \[w(4)=3(4)+7=12+7=19\] \[w(4)=19\]